### v2

# SiamSeg: Self-Training with Contrastive Learning for Unsupervised Domain Adaptation Semantic Segmentation in Remote Sensing

Abstract
Semantic segmentation of remote sensing (RS) images is a challenging yet crucial task. While deep learning, particularly supervised learning with large-scale labeled datasets, has significantly advanced this field, acquiring high-quality labeled data is expensive and time-consuming. Additionally, variations in ground sampling distance, imaging equipment, and geographic differences cause domain shifts between datasets, which limit model performance across domains. Unsupervised domain adaptation (UDA) offers a solution by enabling models to learn from unlabeled target domain data while training on labeled source domain data. Recent self-supervised learning approaches using pseudo-label generation have shown promise in addressing domain discrepancies. Combining source and target images with their true and pseudo labels has proven effective in reducing domain bias. However, the use of pseudo-labeling for RS image segmentation is underexplored. Existing methods often rely on high-confidence pixel points as pseudo-labels, reducing supervision in low-confidence areas. Noise in pseudo-labels further weakens the model’s ability to learn target domain semantics. While some methods assign confidence weights, noisy pseudo-labels remain an issue. To address these limitations, we propose integrating contrastive learning into UDA, enhancing the model’s capacity to capture semantic information by maximizing the similarity between augmented views of the same image. This provides additional supervision to improve performance in the target domain. Extensive experiments on key RS datasets (Potsdam, Vaihingen, LoveDA) demonstrate that our SimSeg method outperforms existing approaches, achieving state-of-the-art results. Visualization and quantitative analyses confirm its superior ability to learn from the target domain. The code is available at https://github.com/woldier/SiamSeg.
Index Terms:
Unsupervised Domain adaptation (UDA), Contrastive Learning, Remote Sensing (RS), Semantic Segmentation.
I Introduction
Remote sensing techniques are widely employed in various visual tasks, including remote sensing scene classification [1] [2] [3] [4], target detection [5] [6] [7], and semantic segmentation [8] [9] [10] [11] [12]. Among these, the semantic segmentation task aims to accurately classify each pixel in remote sensing images for pixel-level target recognition. The extensive application of remote sensing image segmentation in urban planning, flood control, and environmental monitoring [13] has garnered increasing attention from researchers, prompting deeper exploration of the topic.
In recent years, deep learning-based semantic segmentation algorithms have made significant strides, leading to the emergence of many top-performing models, such as Fully Convolutional Neural Networks (FCNs) [14] [15] [16] and Transformer-based models [17] [18] [19]. However, the effectiveness of these methods heavily depends on the distributional similarity between the training and test data. When a domain shift occurs between different datasets, model performance significantly deteriorates. In practice, this domain shift problem is particularly pronounced due to the diversity of geographic regions, imaging conditions, and equipment used in remote sensing datasets, resulting in insufficient generalization capability of existing methods.
To address the domain bias problem and establish effective associations between source and target domains, cross-domain semantic segmentation of remote sensing images has emerged as a significant research direction. Unsupervised Domain Adaptation (UDA), a subset of transfer learning, aims to tackle the generalization challenge when the source domain has labeled data while the target domain contains only unlabeled data. Existing UDA methods can be broadly categorized into three groups: adversarial learning-based methods, generative model-based methods, and self-supervised learning (ST)-based methods [20].
Adversarial learning-based approaches assist segmentation networks in minimizing the differences between the distributions of feature maps in the source and target domains by introducing a discriminator [21] [22]. Specifically, the segmentation network first generates feature maps from input images in both domains and makes segmentation predictions. Meanwhile, the discriminator receives these feature maps as input and determines whether they originate from the source domain or the target domain. The segmentation network then aims to deceive the discriminator by adjusting the generated target domain feature maps to align closely with the source domain feature maps. Although adversarial learning methods have achieved some success in narrowing the distribution gap between source and target domains, they have significant drawbacks. First, the training process of adversarial learning is often unstable, particularly when dealing with complex feature spaces, making it prone to issues such as gradient vanishing or explosion. Second, as adversarial learning essentially involves a minimax game, the model may collapse during training; that is, the segmentation network may fail to generate diverse features and instead converge to a single solution, resulting in insufficient generalization ability. These challenges render adversarial learning methods difficult to apply in practical scenarios, particularly in cross-domain segmentation of remote sensing images.
In contrast to adversarial learning-based approaches, generative domain adaptation methods achieve visual similarity between source and target domain images at the pixel level using image-to-image translation or style migration techniques [23]. Specifically, these methods pass source domain images through a style translation network to generate images visually similar to the target domain. Subsequently, these style-migrated images are used to train a segmentation network alongside the labeled data from the source domain, improving performance on the target domain. Although generative methods minimize differences between source and target domains at the pixel level, they also have notable limitations. First, these methods typically require additional generative or style transformation networks, increasing model complexity and computational overhead. Second, the performance of the generative network directly impacts the final performance of the segmentation network. A style transformation network that fails to accurately capture the visual differences between the source and target domains or introduces noise and artifacts during the transformation process may hinder the segmentation network’s effectiveness on the target domain. Thus, this approach heavily relies on the conversion ability of the style transformation network, posing challenges for robustness and generalization in complex cross-domain scenarios.
Unlike adversarial learning and generative domain adaptation methods, self-training (ST) methods [13] do not rely on additional networks. The ST strategy facilitates cross-domain knowledge transfer by generating pseudo-labels [24] [25] [26] [27] [28]. The core idea is to follow the ”from easy to difficult” principle, prioritizing pixels with higher prediction confidence in the target domain as pseudo-label points. Labeled source domain data and pseudo-labeled target domain data are then input into the network for training, gradually improving model performance in the target domain. In remote sensing image segmentation, ST methods are relatively underexplored; this paper aims to apply pseudo-label-based ST methods to remote sensing semantic segmentation tasks. The advantages of ST methods include not needing additional generative networks and enhancing the target domain’s generalization ability through iterative optimization of pseudo-labels. Through this process, the pseudo-labels increasingly approximate the true distribution of the target domain, effectively improving segmentation network performance. However, challenges remain in the current ST approach. Most existing pseudo-label generation methods select only high-confidence pixel points for training, which, while ensuring the reliability of pseudo-labels to some extent, neglects low-confidence regions. This leads to insufficient supervised signals in these areas, preventing the network from fully learning the associated semantic information. Additionally, while some pseudo-labeling methods assign confidence weights to all pixels, they may introduce noise during pseudo-label generation, leading to misclassification of certain pixels. This noise weakens the supervised signals and can cause the network to learn inaccurate semantic information in the target domain. Consequently, although pseudo-labeling-based approaches partially address the domain adaptation challenge, they struggle with low-confidence regions and noisy labels, highlighting the urgent need for effective utilization of unlabeled target domain data.
Figure 1: illustrates the main challenges in the task of semantic segmentation of remote sensing images across domains. These challenges include the problem of domain bias due to ground sampling distances, sensor differences, and variations in geographic landscapes, which affect the model’s ability to generalize across different datasets. From top to bottom, the first section shows the impact of different ground sampling distances from Potsdam IRRG to Vaihingen IRRG, the second section shows the remote sensing sensor variation from Potsdam RGB to Vaihingen IRRG, and the third section compares the geographic landscape differences between rural and urban areas in LoveDA. Details of the dataset will be discussed at IV-A. Dashed boxes in each case mark the differences in the same category across domains, highlighting the impact of these issues on semantic segmentation performance. Understanding these domain bias issues is crucial for improving the accuracy and robustness of semantic segmentation of remotely sensed images and points the way for subsequent research.
More importantly, although many classical unsupervised domain adaptation (UDA) methods have been successfully applied to natural scenes, the domain shift problem in remote sensing images is more complex, stemming from factors such as ground sampling distances, sensor type discrepancies, and geographic landscape variations. As illustrated in Fig. 1, this complexity significantly degrades the performance of algorithms that work well in natural scenes when directly applied to remote sensing data. Specifically, self-supervised methods relying on pseudo-label generation show diminished richness in learned semantic features under complex domain shifts, resulting in lower confidence levels for generated pseudo-labels. The increased noise in these pseudo-labels raises the likelihood of pixel misclassification, further impairing the network’s ability to semantically characterize the target domain.
The rise of contrast learning in computer vision demonstrates its powerful capability to capture semantic information in images without relying on labeled data, resulting in enhanced feature representation. Based on this observation, this paper proposes SiamSeg, which introduces contrast learning for the first time to the unsupervised domain adaptation task of semantic segmentation in remote sensing images. Leveraging the robust feature representation capability of contrast learning, SiamSeg effectively addresses the insufficient semantic information learning caused by the weak supervisory signals of pseudo-labels in the target domain, significantly enhancing segmentation network performance. To the best of our knowledge, SiamSeg is the first UDA approach to exploit contrast learning for facilitating context relations learning in the target domain.
  1. 1.
    Given the limited exploration of pseudo-labeled self-supervised learning in remote sensing UDA segmentation, this study focuses on the ST approach for UDA.
  2. 2.
    This paper presents the first application of contrast learning to an UDA task. Positive sample pairs are generated through different data augmentations of the same image, and the model is optimized using a Siamese network.
  3. 3.
    A novel loss function, based on contrastive learning, is proposed that incorporates contrast learning loss to enhance the model’s learning effectiveness.
  4. 4.
    Comprehensive experiments are conducted, demonstrating that the method achieves new state-of-the-art performance across various datasets
II Related Work
II-A Unsupervised cross-domain adaptation for semantic segmentation
Adversarial learning is a prevalent approach among various effective methods. Tsai et al. [29] argued that there is a high degree of similarity between the source and target domains in terms of semantic layout, leading them to construct a multi-level adversarial network that exploits structural consistency in the cross-domain output space. Conversely, Vu et al. aimed to minimize the difference between the entropy distributions of the source and target domains by introducing a discriminator [30]. Cai et al. proposed a bidirectional adversarial learning framework to maintain semantic consistency in the segmentation of remote sensing images [31].
However, the sensitivity of the discriminator network to hyperparameters presents challenges in learning similar feature distributions across different domains. Additionally, adversarial learning may induce unstable training, further complicating the model’s learning process.
Another typical non-adversarial unsupervised domain adaptation paradigm is self-training, which has gained significant attention in cross-domain semantic segmentation in recent years. Zou et al. [26] first introduced a self-supervised learning method for unsupervised domain-adaptive semantic segmentation, wherein pseudo-labels are treated as discrete latent variables and computed by minimizing a uniform training objective. Tranheden et al., Zhou et al., Hoyer et al., and Chen et al. enhanced domain adaptation by generating trustworthy, consistent, and category-balanced pseudo-labels [27] [32] [33] [34].
However, there has been relatively limited exploration of self-supervised learning methods in the realm of semantic segmentation for remote sensing images. Therefore, this paper seeks to apply self-supervised learning to remote sensing image segmentation tasks. Despite the success of self-supervised learning methods in various applications, several challenges remain. For instance, current pseudo-label generation methods typically select only high-confidence pixels, leading to weakened supervised signals in low-confidence regions, which adversely affects the network’s ability to learn semantic information about the target domain. Additionally, self-supervised learning may introduce noise during the generation of pseudo-labels, resulting in some pixels being incorrectly assigned to erroneous pseudo-label categories. This further limits the model’s performance. In light of these challenges, integrating the advantages of contrastive learning may offer new perspectives for enhancing the application of self-supervised learning in remote sensing image segmentation.
II-B Contrastive Learning
The core principle of contrastive learning is to generate pairs of images (view pairs or positive sample pairs) that share the same potential significance [35]. The optimization objective of contrastive learning is to encourage the model to learn similar embeddings for positive sample pairs while effectively distinguishing irrelevant sample pairs (negative sample pairs). This approach has gained prominence in unsupervised self-training representation learning [36] [37] [38]. The concept of simple and effective contrastive learning was further advanced through the introduction of the Siamese network [39] [40] [41] [42] [43] [44].
In practice, the performance of contrastive learning methods is significantly enhanced, partly due to the utilization of a large number of negative samples, which can be stored in a memory bank [36]. For instance, the MoCo method [41] maintains a queue of negative samples and employs a momentum encoder to improve the consistency of this queue. In contrast, the SimCLR method directly utilizes negative samples present in the current batch, though it typically requires a larger batch size to function effectively. The SimSiam method [44] achieves effective feature learning by simplifying the design. Unlike other contrastive learning methods, SimSiam does not rely on negative samples but instead builds “pairs of positive samples” for training. During training, the model processes different augmented versions of the same image and extracts features through a two-branch network. The two branches propagate forward separately, with a key component being the stop-gradient operation, which ensures that the parameters of one branch are not updated during backpropagation. This design facilitates the network’s learning capacity. SimSiam’s architecture allows it to perform well on small datasets, offering a new perspective on contrastive learning methods.
In the context of remote sensing (RS) images, the richer image features can lead to insufficient supervisory signals when using self-supervised learning with pseudo-labels. This limitation may hinder the effective learning of image features, adversely affecting the segmentation performance of the model. Therefore, enhancing the model’s representation learning ability through contrastive learning is essential, as it not only improves the accuracy of feature learning but also enhances the model’s performance in complex scenes.
III Method
III-A Preliminary
In Unsupervised Domain Adaptation (UDA) for Remote Sensing (RS), we define two sets of images collected from different satellites or locations as distinct domains. To simplify the problem, we assume that the source domain and target domain images share the same pixel resolution, denoted as $H\times W\times 3$, where $H$ and $W$ represent the image height and width, respectively, and $3$ indicates the number of channels. Additionally, the two domains maintain consistency in the number of classes.
Let $x^{(i)}_{S}$ be the image and $y^{(i)}_{S}$ its corresponding label, with the source domain defined as $D_{S}=\left\{(x_{S}^{(i)},y_{S}^{(i)})\mid x_{S}^{(i)}\in\mathbb{R}^{H\times W\times 3},y_{S}^{(i)}\in\mathbb{R}^{H\times W\times C}\right\}_{i=1}^{N_{S}}$ , where $C$ is the number of classes. The target domain is defined as $D_{T}=\left\{x_{T}^{(i)}\mid x_{T}^{(i)}\in\mathbb{R}^{H\times W\times 3}\right\}_{i=1}^{N_{T}}$ , where only the images $x_{T}^{(i)}$ are accessible, while the labels $y_{T}^{(i)}$ remain unavailable. The subscripts $S$ and $T$ denote the source and target domains, respectively, and $N_{S}$ and $N_{T}$ indicate the number of samples in the source and target domains. It is important to note that during training, the labels $y_{T}^{(i)}$ from the target domain are inaccessible, being utilized only during the testing phase. The representation of the source domain label $y_{S}$ at the spatial position $(h,w)$ is denoted as a length $C$ one-hot encoding, represented as $y^{(i,h,w)}_{S}$, where $h\in[1,\dots,H]$ and $w\in[1,\dots,W]$.
The goal of UDA is to leverage the source domain images $x_{S}^{(i)}$ and their labels $y_{S}^{(i)}$ to train the model, enabling it to perform well on target domain images $x_{T}^{(i)}$ without access to the target domain labels $y_{T}^{(i)}$. If we solely rely on cross-entropy loss in the source domain for training the network $g_{\theta}$, it can be expressed as follows:
$L^{(i)}_{S}=-\sum_{h=1}^{H}\sum_{w=1}^{W}\sum_{c=1}^{C}y^{(i,h,w,c)}_{S}\cdot log(g_{\theta}(x^{(i)}_{S})^{(h,w,c)}).$ (1)
In Equation (1), $g_{\theta}(x^{(i)}_{S})$ represents the predicted outcomes for each pixel in the source domain image $x^{(i)}_{S}$, providing normalized probabilities for each class at the spatial position $(h,w)$. However, due to the domain gap, relying solely on the source domain for training typically results in poor performance on the target domain, as the network struggles to generalize to target domain samples.
III-B Self-Training (ST) for UDA
To effectively address the domain gap, numerous methods have been proposed, broadly categorized into adversarial training [45] [29] [30] [31] and Self-Training (ST) [26] [27] [46]. Due to the instability of adversarial training, self-training has demonstrated superior performance across various tasks [27] [46]. Particularly in the field of remote sensing, research on ST methods is relatively scarce, which is why this study opts for a self-training approach.
To transfer knowledge from the source domain to the target domain, the ST method employs a teacher network $t_{\theta}$ to generate corresponding pseudo-labels for the target domain images (a detailed discussion on the teacher network will be presented in III-C). Mathematically, this is expressed in Equation (2):
$p_{T}^{(i,h,w,c)}=\left[c=arg\max_{c^{\prime}}t_{\theta}(x^{(i)}_{T})^{(h,w,c)}\right].$ (2)
where []delimited-[][$\cdot$][ ⋅ ] denotes the Iverson bracket. Here, $t_{\theta}(x_{T}^{(i)})$ indicates the class predictions for each pixel in the target domain image $x^{i}_{T}$. It is important to note that gradients are not backpropagated through the teacher network. Since we cannot ascertain the correctness of the generated pseudo-labels, it is necessary to evaluate the quality or confidence of the predictions for the pixels in the pseudo-labels. Only those pixels with class confidence exceeding a threshold $\tau$ will be used for training. The mathematical formulation for assessing the quality or confidence of the pixel at position $(h,w)$ is given by equation (3):
$q_{T}^{(i)}=\frac{{\textstyle\sum_{h=1}^{H}}{\textstyle\sum_{w=1}^{W}}[\max(t_{\theta}(x_{T}^{(i)})^{(h,w)})]>\tau}{H\cdot W}.$ (3)
The meaning of []delimited-[][$\cdot$][ ⋅ ] is consistent with that in Equation (2). We choose $t_{\theta}$ to generate pseudo-labels because it is more stable than the student network $g_{\theta}$. Higher thresholds generally correspond to higher-quality pseudo-labels; however, this may result in fewer pseudo-label pixels being utilized for training during the early phases. As the training iterations increase, confidence gradually rises, leading to an increase in the number of pseudo-labels employed.
In addition to training on labeled data in the source domain using Equation (1), the pseudo-labels $p_{T}^{(i)}$ and their corresponding quality estimates $q_{T}^{(i)}$ from the target domain will also be incorporated into the training of the student network $g_{\theta}$. The loss for training in the target domain can be mathematically represented as Equation (4):
$L^{(i)}_{T}=-\sum_{h=1}^{H}\sum_{w=1}^{W}\sum_{c=1}^{C}q_{T}^{(i)}\cdot p^{(i,h,w,c)}_{T}\cdot log(g_{\theta}(x^{(i)}_{T})^{(h,w,c)}).$ (4)
Here, $g_{\theta}(x^{(i)}_{T})$ denotes the predicted outcomes for each pixel in the target domain image $x^{(i)}_{T}$, providing normalized probability predictions. Through this approach, the model can effectively utilize pseudo-label information to enhance its performance in the target domain.
III-C Pseudo-Label Generation
The generation of pseudo-labels $p_{T}^{(i)}$ can be conducted through online [47] [27] or offline [25] [48] [26] methods. In online self-training, the teacher network $t_{\theta}$ updates in tandem with $g_{\theta}$ during the training process. This procedure is relatively straightforward; however, it is crucial to note that the structures of the teacher network $t_{\theta}$ and the student network $g_{\theta}$ are identical. Moreover, due to the use of the stop-gradient operation, the teacher network does not backpropagate gradients during training. Typically, to mitigate the parameter perturbations caused by the domain gap during the training process, the weights of the teacher network are updated after each training iteration using the exponentially moving average (EMA) method [49], thereby enhancing the stability of pseudo-label generation. This can be mathematically expressed as Equation (5):
$S_{t+1}(t_{\theta})\leftarrow\alpha\cdot S_{t}(t_{\theta})+(1-\alpha)\cdot S_{t}(g_{\theta}).$ (5)
where $S_{t}(\cdot)$ denotes the weights of the model at training step $t$, and the hyperparameter $\alpha\in[0,1]$ indicates the importance of the current state $S(t_{\theta})$. This weight update strategy not only enhances the accuracy of pseudo-label generation but also alleviates the adverse effects of domain differences, ultimately improving the model’s generalization ability in the target domain.
While pseudo-labeling methods have made significant progress in addressing domain adaptation, they still face challenges. In particular, noisy pseudo-labels can arise from low-confidence regions, leading to pixel misclassification and weakening the supervised signals needed for effective learning in the target domain. As domain shifts in remote sensing images become more complex—due to factors like sensor discrepancies and variations in geographic landscapes—the accuracy of pseudo-labels tends to degrade. This increased noise introduces additional difficulty in training the model to semantically characterize the target domain, limiting the overall effectiveness of these methods.
III-D Contrastive Learning
To overcome the limitations of pseudo-labeling, we introduce a contrastive learning. The contrastive learning module generates two distinct views $x_{T}^{(i)}$ through two random augmentations. These views are processed by an encoder network $E$ , consisting of a backbone $f$ (e.g., MIT [18]) and a Multi-Layer Perceptron (MLP) projection head [42]. The weights of the encoder network $E$ remain the same while processing both views. We denote the MLP prediction head as $h$; the output of one view through the encoder network $E$ is transformed to match the representation of the other view. The outputs of the MLP prediction head and the MLP projection head can be expressed as $p_{1}^{(i)}=h(E(x_{T_{1}}^{(i)}))$ and $z_{1}^{(i)}=E(x_{T_{2}}^{(i)})$. Our objective is to optimize the negative cosine similarity between these two vectors, mathematically represented as Equation (6):
$D(p_{1}^{(i)},z_{2}^{(i)})=-\frac{p_{1}^{(i)}}{\left\|p_{1}^{(i)}\right\|_{2}}\cdot\frac{z_{2}^{(i)}}{\left\|z_{2}^{(i)}\right\|_{2}}.$ (6)
where $\left\|\cdot\right\|_{2}$ denotes $l_{2}$ normalization. Thus, we define the contrastive learning loss, in Equation (7), as follows:
$L_{CLR}^{i}=\frac{1}{2}\cdot D(p_{1}^{(i)},stop\_grad(z_{2}^{(i)}))+\frac{1}{2}\cdot D(p_{2}^{(i)},stop\_grad(z_{1}^{(i)})).$ (7)
In the loss function, the stop-gradient operation is a crucial step, treating the variables in $stop\_grad(\cdot)$ as constants. This operation effectively prevents representation collapse [44]. Specifically, in the first term of Equation (7), the encoder processing $x_{T_{2}}^{(i)}$ does not receive gradients from $z_{2}^{(i)}$, while in the second term, the encoder processing $x_{T_{1}}^{(i)}$ does receive gradients from $p_{2}^{(i)}$. This design effectively enhances the model’s representation learning capabilities, improving its performance in the target domain.
III-E Proposed UDA Losses
In a given training step, we acquire the $m$-th image and its corresponding label from the source domain, represented as $x_{S}^{(m)}$ and $y_{S}^{m}$. Simultaneously, we obtain the $m$-th image from the target domain and its pseudo-label generated by the teacher network $t_{\theta}$, denoted as $x_{T}^{(n)}$ and $p_{T}^{(n)}$. The total loss function we mathematically define is as follows:
$L_{total}^{m;n}=L_{S}^{(m)}+\beta\cdot L_{T}^{(n)}+\gamma\cdot L_{CLR}^{(n)}.$ (8)
where $\beta$ and $\gamma$ are balancing factors used to weigh the different losses. Through Equation (8), we can comprehensively consider the losses from both the source and target domains, facilitating more effective model training.
Specifically, the loss $L_{S}^{(m)}$ guides the model’s performance on source domain data, while $L_{T}^{(n)}$ leverages pseudo-labels to drive learning in the target domain, further enhancing the model’s generalization ability in this domain. Finally, $L_{CLR}^{(n)}$ introduces richer feature representations through contrastive learning, enabling the model to better adapt to potential differences between the source and target domains. By integrating these losses, we can effectively reduce the domain gap between the source and target domains, thereby improving the model’s performance on the unlabeled target domain data.
III-F SiamSeg Network Architecture
III-F1 overall
Figure 2: The SiamSeg network $g_{\theta}$ is designed for image segmentation and comprises a feature extraction backbone $f$ and a decoding head. An Exponential Moving Average (EMA) teacher network $t_{\theta}$ enhances stability without backpropagating gradients, relying on the EMA strategy of $g_{\theta}$. During training, the source image $x_{S}^{(m)}$ and label $y_{S}^{(m)}$ train $g_{\theta}$, while the target image $x_{T}^{(n)}$ generates pseudo-labels $p_{T}^{(n)}$ through $t_{\theta}$, facilitating self-training. Data augmentation creates two views for contrastive learning, enriching training signals.
As illustrated in Figure 2, the network $g_{\theta}$ is a model specifically designed for image segmentation tasks. Its architecture comprises a feature extraction backbone $f$ and a decoding head responsible for segmentation predictions. To enhance the model’s stability and performance, we introduce an Exponential Moving Average (EMA) teacher network $t_{\theta}$, which has an identical architecture to $g_{\theta}$ but does not backpropagate gradients during training. This design ensures that the weight updates of $t_{\theta}$ rely solely on the exponential moving average strategy of $g_{\theta}$, effectively reducing fluctuations during the training process.
In the contrastive learning module, the network structure similarly includes the backbone $f$, a projection Multi-Layer Perceptron (MLP) head, and a prediction MLP head. Notably, the backbone $f$ shares weights with the backbone of the segmentation network $g_{\theta}$ to enhance the consistency and effectiveness of feature representations.
During training, as indicated by the blue arrows in the figure, the source image $x_{S}^{(m)}$ and its corresponding source label $y_{S}^{(m)}$ are utilized to initially train the network $g_{\theta}$ through supervised learning. In the step denoted by the black arrows, the target image $x_{T}^{(n)}$ is processed by the EMA teacher network $t_{\theta}$, generating pseudo-labels $p_{T}^{(n)}$ to replace the inaccessible target labels $y_{T}^{(n)}$ during training. This process not only reduces the reliance on true labels but also enhances the model’s adaptability in the target domain. After generating pseudo-labels, the target image $x_{T}^{(n)}$ and pseudo-label $p_{T}^{(n)}$ are combined for Self-Training. It is important to emphasize that during the pseudo-label generation process, the teacher network $t_{\theta}$ does not participate in backpropagation due to the introduction of the stop-gradient operation.
Furthermore, by applying data augmentation to the target image $x_{T}^{(n)}$, we generate two distinct views, which together form positive sample pairs and provide rich training signals, with contrastive learning, for model. The following section III-F2 will explore the specific architecture and implementation details of the contrastive learning network, revealing its significant role in self-supervised learning.
III-F2 detail of contrastive learning
Figure 3: Siamese Network Architecture for Contrastive Learning. This figure illustrates the architecture of the Siamese network used for contrastive learning. The network consists of two identical sub-networks that share the same model weights, ensuring consistency in feature extraction. Each input image $x_{T}^{(n)}$ undergoes various data augmentation techniques (eg., Resize Crop, Color Jitter, Gray Scale, Gaussian Blur, Filp), generating two distinct views: $x_{T1}^{(n)}$ and $x_{T2}^{(n)}$. The first view is processed through an encoder and a prediction MLP head, producing output $\mathbf{p}_{1}^{(n)}$. In contrast, the second view is processed only through the encoder, yielding output $\mathbf{z}_{2}^{(n)}$ while employing a stop-gradient operation to optimize the network. The diagram highlights the symmetrical workflows of both views and the optimization process aimed at maximizing the similarity between $\mathbf{p}_{1}^{(n)}$ and $\mathbf{z}_{2}^{(n)}$, thereby enhancing feature representation for effective contrastive learning.
The implementation of contrastive learning utilizes the Siamese network architecture, a simple yet effective strategy [39] [41] [44]. As shown in Fig. 3, the Siamese network comprises two identical sub-networks that share the same model weights, ensuring consistency during feature extraction and facilitating effective contrastive learning. Specifically, an input image $x_{T}^{(n)}$ undergoes data augmentation (eg., Resize Crop, Color Jitter, Gray Scale, Gaussian Blur, Filp), generating two distinct views, $x_{T1}^{(n)}$ and $x_{T2}^{(n)}$. These views are processed by the contrastive networks. Despite sharing parameters, their outputs differ slightly. In the upper workflow, the view $x_{T1}^{(n)}$ is first encoded by the encoder $E$ and then processed through the prediction MLP head $h$ to yield output $\mathbf{p}_{1}^{(n)}$. In contrast, the view $x_{T2}^{(n)}$ is also passed through the encoder to produce output $\mathbf{z}_{2}^{(n)}$ but does not go through the prediction MLP head; instead, it employs a stop-gradient operation. To maximize the similarity between outputs $\mathbf{p}_{1}^{(n)}$ and $\mathbf{z}_{2}^{(n)}$, we optimize using Equation 7. This process encourages the model to fully exploit the similarities between positive sample pairs, thereby enhancing the effectiveness and robustness of feature representation. The lower workflow mirrors the upper one, ensuring the unity and consistency of the overall contrastive learning process. Through this structured design, the Siamese network effectively performs feature learning, providing strong representational capabilities for our RS tasks.
IV EXPERIMENTAL RESULTS AND ANALYSIS
IV-A Datasets
To evaluate the proposed method’s performance in cross-domain remote sensing (RS) image segmentation tasks, we selected three benchmark datasets: Potsdam, Vaihingen, and LoveDA.
IV-A1 Potsdam and Vaihingen datasets
The Potsdam and Vaihingen datasets are part of the ISPRS 2D semantic segmentation benchmarks [50]. The Potsdam (POT) dataset comprises $38$ remote sensing images with a resolution of $6000\times 6000$ pixels and a ground sampling distance of $5$ meters. It features three different image modalities: IRRG, RGB, and RGBIR, where the first two modalities have three channels, and the last one has four. In this study, we primarily utilize the first two modalities. The Vaihingen (VAI) dataset contains $33$ remote sensing images with resolutions ranging from $1996\times 1996$ to $3816\times 2550$ pixels, with a ground sampling distance of $9$ centimeters. This dataset includes only one image modality (IRRG). Both datasets share six common classes: impervious surfaces, buildings, low vegetation, trees, cars, and clutter/background, where the clutter/background class includes water boundaries and other irrelevant objects for semantic segmentation. In our study, we utilize the IRRG and RGB image modalities. Notably, the visual appearance of some objects may differ in the IRRG modality compared to conventional images; for instance, trees and low vegetation appear red.
To reduce computational overhead, we cropped the images to a smaller size of $512\times 512$ pixels. For the POT and VAI datasets, we used cropping strides of $512$ and $256$, resulting in $4598$ and $1696$ images, respectively. Subsequently, we split the POT and VAI datasets into training and testing sets, following previous wokrs [13] [51]. In the POT dataset, the number of images in the training and testing sets is $2904$ and $1694$, respectively, while in the VAI dataset, these numbers are $1296$ and $440$.
Due to the urban nature of Potsdam, it contains a higher number of vehicles, whereas Vaihingen includes more residential and agricultural areas. Although both datasets have the same task objectives and label space, they exhibit significant differences in feature distribution, making them well-suited for evaluating the performance of unsupervised domain adaptation (UDA) methods. Accordingly, we established four cross-domain remote sensing semantic segmentation tasks:
  • Potsdam IR-R-G to Vaihingen IR-R-G (POT IRRG → VAI IRRG).
  • Vaihingen IR-R-G to Potsdam IR-R-G (VAI IRRG → POT IRRG).
  • Potsdam R-G-B to Vaihingen IR-R-G (POT RGB → VAI IRRG).
  • Vaihingen IR-R-G to Potsdam R-G-B (VAI IRRG → POT RGB).
IV-A2 LoveDA dataset
The LoveDA dataset was recently proposed to address semantic segmentation and domain adaptation challenges in remote sensing. It consists of $5987$ high-resolution ($1024\times 1024$) RS images sourced from Nanjing, Guangzhou, and Wuhan [52]. The LoveDA dataset contains two distinct domains: urban and rural, aimed at challenging the model’s generalization capability between different geographical elements.
Within the dataset, there are $1883$ urban images, which are further divided into $1156$ training samples and $677$ validation samples as mentioned in [52]. In the rural domain, there are $2358$ images, with $1366$ for training and $992$ for validation. For the LoveDA dataset, we designed a cross-domain remote sensing semantic segmentation task:
  • Rural to Urban (Rural → Urban).
Through these five different tasks, we aim to assess the proposed method’s adaptability and performance across various geographical environments, providing new perspectives and data support for cross-domain semantic segmentation research.
IV-B Evaluation Metrics
In this study, we adopt the F1-score and Intersection over Union (IoU) as evaluation metrics, following previous methods in the field of remote sensing (RS) semantic segmentation domain adaptation (UDA). The specific calculation formulas are as follows:
$\text{IoU}=\frac{TP}{TP+FP+FN}.$ (9)
[…]
In equations (9) and (10), $TP$ represents true positives, $FP$ denotes false positives, and $FN$ indicates false negatives. The IoU, also known as the Jaccard index, and the F1-score, referred to as the Dice coefficient, effectively reflect the accuracy and reliability of model performance in semantic segmentation tasks. These metrics are particularly valuable when addressing imbalanced datasets, as the F1-score provides a more comprehensive evaluation of performance. By utilizing these evaluation metrics, we can systematically compare the performance of different methods in cross-domain RS image segmentation tasks.
IV-C Implementation Details
IV-C1 Augmentation
This study is implemented using the PyTorch framework [53] and MMSegmentation [54]. We utilize the data preprocessing pipeline provided by MMSegmentation, which includes operations such as random image resizing, cropping, and flipping. In addition, inspired by the DACS approach [27], we incorporate color jitter, Gaussian blur, and ClassMix [55] to enhance the dataset and improve feature robustness. During the contrastive learning stage, we generate two different views of the image by applying augmentations such as cropping, color jitter, Gaussian blur, grayscale, and flipping.
IV-C2 Network Architecture Details
Considering the outstanding performance of the Transformer-based Segformer [18] in semantic segmentation tasks, we chose Mix Vision Transformers (MiT) [18] as the backbone network (denoted as backbone $f$). This model was pre-trained on ImageNet [56], eliminating the need to use convolutional neural network (CNN)-based architectures like DeepLabV3+ [57]. In semantic segmentation, capturing both global and local features is critical, and feature fusion strategies are often used to improve segmentation performance. For this reason, we adopt the context-aware multi-scale feature fusion decoder head designed by Hoyer et al. [33], given its superior performance in UDA-based semantic segmentation. The feature pyramid generated by the backbone $f$ has channel sizes of [64, 128, 320, 512]. The decoder head projects all features to 256 channels while maintaining a resolution of $\frac{H}{4}\times\frac{W}{4}$ for further processing and segmentation tasks.
IV-C3 Training
During training, the AdamW optimizer is applied to network $g_{\theta}$, with hyperparameters set as betas=(0.9, 0.999) and a weight decay of 0.01. The learning rate for the backbone $f$ is set to $6\times 10^{-4}$, while the decoder head, projection MLP head, and prediction MLP head have learning rates of $6\times 10^{-5}$. For learning rate scheduling, a linear warm-up strategy is employed for the first 1,500 iterations, followed by linear decay with a decay rate of 0.01. The decay coefficient $\alpha$ for the exponential moving average teacher network $t_{\theta}$ is set to 0.99.
In the pseudo-label generation phase, the temperature parameter $\tau$ in Equation 2 is set to 0.999. The total number of training iterations is 40,000, with a batch size of 12, containing 6 source domain and 6 target domain images. During the preprocessing of the Potsdam (POT) and Vaihingen (VAI) datasets, image scaling is set to $img\_scale=(512,512)$, and the aspect ratio range is set to $ratio\_range=(1.0,2.0)$. Crops are sized at $512\times 512$, with $cat\_max\_ratio=0.75$ and a flipping probability of 0.5. Normalization is performed using the mean values [123.675, 116.28, 103.53] and standard deviations [58.395, 57.12, 57.375].
For the LoveDA dataset, we retain the settings from POT and VAI but adjust certain parameters. Specifically, image scaling is set to $img\_scale=(1024,1024)$ and the aspect ratio range to $ratio\_range=(0.5,1.5)$. During the contrastive learning process, data augmentations (Sim. Aug.) use a crop size of $size=(512,512)$ and scale range $scale=(0.6,1.0)$. Color jitter parameters are set to a brightness, contrast, saturation, and hue of 0.25 each, with a random application probability of 0.6. Grayscale has a probability of 0.2, Gaussian blur has a probability of 0.5, and both horizontal and vertical flips are applied with a probability of 0.5. These settings aim to enhance model generalization through data augmentation, improving performance in cross-domain remote sensing image segmentation tasks.
IV-C4 Testing and Validation
For the testing and validation stages, we use MultiScaleFlipAug [54] to evaluate semantic segmentation performance, applying multi-scale augmentation and flipping during testing. For the POT and VAI datasets, since the image size is set to $512\times 512$ after dataset preparation, we set $img\_scale=(512,512)$ and use an image scaling ratio of $img\_ratios=[1.0]$. For the LoveDA dataset, where the original image size is $1024\times 1024$, the image scaling is set to $img\_scale=(1024,1024)$ with scaling ratios $img\_ratios=[0.75,1.0,1.25,1.5]$.
All experiments are conducted within the MMSegmentation framework to ensure consistency and reproducibility. Moreover, all model training is performed using Nvidia A100 GPU×$\times 4$, enhancing training efficiency and performance. These settings provide a solid foundation for the subsequent experimental results, ensuring the validity and stability of the model across different datasets.
IV-D Experimental Results
IV-D1 Quantitative Results
Cross-domain RS Image Semantic Segmentation on POT and VAI
For the Potsdam (POT) and Vaihingen (VAI) datasets, as described in Section IV-A, we established four sets of cross-domain remote sensing (RS) semantic segmentation tasks. In this subsection, we validate the effectiveness of the proposed SimSeg method through a series of comprehensive experimental results.
Since the backbone network used in this study is Mix Transformers (MiT) [18], Segformer is selected as the baseline for comparison. The Segformer model was trained solely on the source domain and tested directly on the target domain. Additionally, we evaluate multiple comparison methods, including AdaptSegNet [29], ProDA [46], and several RS-specific segmentation methods, such as DualGAN [58], CIA-UDA [59], and DNT [60]. These comparisons allow us to demonstrate not only the superior performance of SimSeg in cross-domain RS image semantic segmentation but also to provide valuable insights for further research in domain adaptation.
POT IRRG → VAI IRRG. In this task, Potsdam IRRG is used as the source domain, while Vaihingen IRRG is the target domain. POT contains 2904 labeled training images, and VAI contains 1296 unlabeled training images. Additionally, 440 labeled test images from VAI are used to evaluate model performance. The results are shown in Table I.
Even when trained only on the source domain, Segformer outperforms several UDA methods based on the DeeplabV3+ [57] series, such as AdaptSegNet, ProDA, and DualGAN, in certain categories. The SimSeg method, proposed in this paper, achieves a nearly 20% improvement in mIoU and a 17% increase in mF1, thanks to the combination of pseudo-labeling self-training (ST) and contrastive learning (C.L.).
SimSeg, through the use of pseudo-label ST for UDA, surpasses the performance of adversarial learning methods like CIA-UDA and CNT, highlighting ST as a highly competitive approach. Moreover, compared to SiamSeg without C.L., SiamSeg with C.L. exhibits a further 2.53% and 2.61% improvement in mIoU and mF1, respectively. This demonstrates the effectiveness of contrastive learning in providing additional supervision signals, thereby enabling the backbone network to better learn target domain features.
TABLE I: Cross-domain RS image semantic segmentation results from Potsdam IRRG to Vaihingen IRRG. The best and second-best results are highlighted in bold and underlined, respectively, in each column. The evaluation metrics used are IoU and F1-score, where F1-score is abbreviated as F1. All values are presented as percentages (%), with larger values indicating better performance. The last column provides the average scores across all categories. Note that Segformer was trained only on the source domain and then tested directly on the target domain. ”C.L.” in the table refers to contrastive learning as proposed in this study.
Clutter Car Tree Low Vegetation Building Impervious Surface Overall
Method IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 mIoU mF1
Segformer [18] (src.) 4.22 9.47 31.13 47.89 66.31 78.87 44.47 60.38 75.5 87.95 61.03 76.07 46.11 60.11
AdaptSegNet [29] 4.6 8.76 6.4 11.99 52.65 68.96 28.98 44.91 63.14 77.4 54.39 70.39 35.02 47.05
ProDA [46] 3.99 8.21 39.2 56.52 56.26 72.09 34.49 51.65 71.61 82.95 65.51 76.85 44.68 58.05
DualGAN [58] 29.66 45.65 34.34 51.09 57.66 73.14 38.87 55.97 62.3 76.77 49.41 66.13 45.38 61.43
CIA-UDA[59] 27.8 43.51 52.91 69.21 64.11 78.13 48.03 64.9 75.13 85.8 63.28 77.51 55.21 69.84
DNT[60] 14.77 25.74 53.88 70.03 59.19 74.37 47.51 64.42 80.04 88.91 69.74 82.18 54.19 57.61
SiamSeg w./o. C.L. 18.14 30.71 56.57 72.26 73.95 85.03 62.26 76.74 87.47 93.32 79.74 88.73 63.02 74.46
SiamSeg 27.6 43.26 52.51 68.86 77.69 87.44 65.73 79.32 89 94.18 80.75 89.35 65.55 77.07
POT RGB → VAI IRRG. In this task, we utilized Potsdam RGB as the source domain and Vaihingen IRRG as the target domain. The POT dataset contains 2,904 labeled training images, while the VAI dataset has 1,296 unlabeled training images and 440 labeled testing images for performance evaluation. The results, shown in Table II, exhibit a trend similar to the POT IRRG → VAI IRRG task.
Notably, Segformer, as the baseline method, is trained solely on the source domain (POT RGB) and tested directly on the target domain (VAI IRRG). Although Segformer shows relatively balanced performance across all categories, its generalization capability in cross-domain scenarios is limited, achieving an mIoU of 36.95 and an mF1 of 48.17, indicating performance bottlenecks in domain transfer.
For the adversarial learning-based UDA methods, AdaptSegNet and ProDA display significant shortcomings across multiple categories, particularly with AdaptSegNet performing poorly in the “car” and “low vegetation” categories. Although ProDA shows slight improvements in these categories, its overall mIoU and mF1 scores remain unsatisfactory, suggesting inadequate adaptability for remote sensing image semantic segmentation tasks.
Among UDA methods designed for the remote sensing domain, DualGAN demonstrates noteworthy performance in the “car” and “low vegetation” categories, while DNT showcases strong generalization capability, especially in the “building” and “impervious surface” categories. CIA-UDA exhibits excellent performance in the “car” and “tree” categories and displays robust cross-domain adaptability across several classes.
The SiamSeg method without contrastive learning performs well in the “building” category. However, the version with contrastive learning exhibits significant improvements across multiple categories, particularly in “low vegetation” and “impervious surface,” confirming that contrastive learning provides additional supervisory signals that enhance model performance on the target domain.
TABLE II: Cross-domain RS image semantic segmentation results from Potsdam RGB to Vaihingen IRRG. The best and second-best results are highlighted in bold and underlined, respectively, in each column. The evaluation metrics used are IoU and F1-score, where F1-score is abbreviated as F1. All values are presented as percentages (%), with larger values indicating better performance. The last column provides the average scores across all categories. Note that Segformer was trained only on the source domain and then tested directly on the target domain. ”C.L.” in the table refers to contrastive learning as proposed in this study.
Clutter Car Tree Low Vegetation Building Impervious Surface Overall
Method IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 mIoU mF1
Segformer [18] (src.) 1.43 2.81 37.97 55.04 52.62 68.96 5.18 9.85 73.18 84.51 51.34 67.85 36.95 48.17
AdaptSegNet [29] 2.29 5.81 10.25 18.45 55.51 68.02 12.75 22.61 60.72 75.55 51.26 67.77 31.58 43.05
ProDA [46] 2.39 5.09 31.56 48.16 49.11 65.86 32.44 49.06 68.94 81.89 49.04 66.11 38.91 52.7
DualGAN [58] 3.94 13.88 40.31 57.88 55.82 70.61 27.85 42.17 65.44 83 49.16 61.33 39.93 54.82
CIA-UDA [59] 13.5 23.78 55.58 68.66 63.43 77.62 33.31 49.97 79.71 88.71 62.63 77.02 50.81 64.29
DNT [60] 11.55 20.71 52.64 68.97 58.43 73.76 43.63 61.5 81.09 89.56 67.94 80.91 52.6 65.83
SiamSeg w./o. C.L. 6.66 12.49 51.85 68.29 68.06 80.99 28.39 44.22 83.6 91.07 69.23 81.82 51.3 63.15
SiamSeg 13.23 23.37 51.14 67.67 71.09 83.1 40.1 57.24 81.56 89.85 73.15 84.49 55.04 67.62
VAI IRRG → POT IRRG. In this task, we used Vaihingen IRRG as the source domain and Potsdam IRRG as the target domain. The VAI dataset contains 1,296 labeled training images, while the POT dataset comprises 2,904 unlabeled training images, along with 1,694 labeled POT testing images used for performance evaluation of each model. The experimental results are presented in Table III.
In this task, Segformer serves as the baseline method, trained solely on the source domain VAI IRRG and tested on the target domain POT IRRG. Despite its strong performance in the ”building” category, the overall results remain weak, indicating a significant drop in performance during domain transfer.
AdaptSegNet performs inadequately in the ”car” category, while ProDA shows slight improvement in this category and achieves higher IoU in the ”tree” and ”impervious surface” categories. However, both adversarial learning methods maintain a low to moderate overall mIoU, suggesting limited cross-domain adaptability for this task.
DualGAN exhibits some improvement in the ”car” and ”tree” categories compared to AdaptSegNet and ProDA, yet it lacks competitive generalization capabilities and performs averagely in cross-domain adaptation tasks. In contrast, CIA-UDA excels across multiple categories, particularly in ”car” and ”impervious surface,” demonstrating significant advantages. DNT exhibits stable performance across most categories, showcasing reliable cross-domain generalization capabilities.
SiamSeg, by incorporating contrastive learning, significantly enhances performance across all categories, particularly excelling in ”car,” ”building,” and ”impervious surface.” Its mIoU and mF1 scores are the highest among all methods, confirming the effectiveness of contrastive learning in cross-domain semantic segmentation tasks.
TABLE III: Cross-domain RS image semantic segmentation results from Vaihingen IRRG to Potsdam IRRG. The best and second-best results are highlighted in bold and underlined, respectively, in each column. The evaluation metrics used are IoU and F1-score, where F1-score is abbreviated as F1. All values are presented as percentages (%), with larger values indicating better performance. The last column provides the average scores across all categories. Note that Segformer was trained only on the source domain and then tested directly on the target domain. ”C.L.” in the table refers to contrastive learning as proposed in this study.
Clutter Car Tree Low Vegetation Building Impervious Surface Overall
Method IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 mIoU mF1
Segformer [18] (src.) 1.08 2.56 58.99 73.14 30.07 46.24 51.91 68.82 74.85 87.18 60.63 76.47 46.29 59.08
AdaptSegNet [29] 8.36 15.33 40.95 58.11 22.59 36.79 34.43 64.5 48.01 63.41 49.55 64.64 33.98 49.96
ProDA [46] 10.63 19.21 46.78 63.74 31.59 48.02 40.55 57.71 56.85 72.49 44.7 61.72 38.51 53.82
DualGAN [58] 11.48 20.56 48.49 65.31 34.98 51.82 36.5 53.48 53.37 69.59 51.01 67.53 39.3 54.71
CIA-UDA [59] 10.87 19.61 65.35 79.04 47.74 64.63 54.4 70.47 72.31 83.93 62.74 77.11 52.23 65.8
DNT [60] 11.51 20.65 49.5 66.22 35.46 52.36 37.61 54.67 66.41 79.81 61.91 76.48 43.74 58.36
SiamSeg w./o. C.L. 3.04 5.91 76.29 86.55 58.51 73.82 66.67 80 83.7 91.13 75.87 86.28 60.68 70.62
SiamSeg 5.2 9.89 76.19 86.48 63.22 77.47 67.54 80.62 83.16 90.8 76.3 86.56 61.94 71.97
VAI IRRG → POT RGB. In this task, we selected Vaihingen IRRG as the source domain and Potsdam RGB as the target domain. The Vaihingen dataset (VAI) contains a total of 1,296 labeled training images, while the Potsdam dataset (POT) includes 2,904 unlabeled training images. Additionally, the POT dataset contains 1,694 labeled testing images used to evaluate the cross-domain performance of each model. The experimental results are detailed in Table IV.
Overall, SiamSeg demonstrates superior performance in cross-domain adaptation tasks, particularly excelling in the ”car,” ”tree,” and ”building” categories, achieving the best IoU and F1 scores. The incorporation of contrastive learning effectively enhances the model’s generalization capability in the target domain, significantly improving cross-domain segmentation performance. In contrast, CIA-UDA also shows strong competitiveness, particularly in the ”car” and ”building” categories, achieving results close to those of SiamSeg, indicating its excellent transferability in certain categories.
Conversely, other adversarial learning-based methods, such as ProDA and DualGAN, display relatively weaker overall performance. Although they show improvements in some categories, they fail to consistently enhance the cross-domain segmentation capability across all categories. For example, ProDA performs well in the ”low vegetation” category but demonstrates limited performance in the ”building” and ”impervious surface” categories. Similarly, DualGAN, while achieving robust IoU in the ”tree” and ”building” categories, still underperforms in others, lacking consistent cross-domain adaptability.
The results in Table IV further validate the cross-domain advantages of SiamSeg following the introduction of contrastive learning, particularly in multi-class segmentation tasks, where mIoU and mF1 metrics are significantly improved. Meanwhile, CIA-UDA, as a competitor, also exhibits good cross-domain generalization capabilities, especially in key target categories. In comparison, traditional adversarial learning methods show uneven performance, with notable limitations in cross-domain segmentation effectiveness.
TABLE IV: Cross-domain RS image semantic segmentation results from Vaihingen IRRG to Potsdam RGB. The best and second-best results are highlighted in bold and underlined, respectively, in each column. The evaluation metrics used are IoU and F1-score, where F1-score is abbreviated as F1. All values are presented as percentages (%), with larger values indicating better performance. The last column provides the average scores across all categories. Note that Segformer was trained only on the source domain and then tested directly on the target domain. ”C.L.” in the table refers to contrastive learning as proposed in this study.
Clutter Car Tree Low Vegetation Building Impervious Surface Overall
Method IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 mIoU mF1
Segformer [18] 2.36 4.61 72.16 83.83 5.38 10.21 31.52 48.65 72.61 84.13 62.45 76.89 41.08 51.39
AdaptSegNet [29] 6.11 11.5 42.31 55.95 30.71 45.51 15.1 25.81 54.25 70.31 37.66 59.55 31.02 44.75
ProDA [46] 11.13 20.51 41.21 59.27 30.56 46.91 35.84 52.75 46.37 63.06 44.77 62.03 34.98 50.76
DualGAN [58] 13.56 23.84 39.71 56.84 25.8 40.97 41.73 58.87 59.01 74.22 45.96 62.97 37.63 52.95
CIA-UDA [59] 9.2 16.86 63.36 77.57 44.9 61.97 43.96 61.07 70.48 82.68 53.39 69.61 47.55 61.63
DNT [60] 8.43 15.55 46.78 63.74 36.56 53.55 30.59 46.85 69.95 82.32 56.41 72.13 41.45 55.69
SiamSeg w./o. C.L. 6.99 13.07 67.75 80.77 55.82 71.65 51.72 68.17 79.03 88.29 65.46 79.12 54.46 66.85
SiamSeg 6.69 12.55 66.76 80.07 57.96 73.39 53.34 69.57 81.52 89.82 68.15 81.06 55.74 67.74
Cross-domain RS image semantic segmentation on LoveDA Rural to Urban
In Section IV-A, we established a cross-domain remote sensing image semantic segmentation task using the LoveDA dataset to validate the effectiveness of the proposed SimSeg method. In this section, we selected several representative comparative methods, including AdaptSegNet [29], FADA [61], CLAN [62], PyCDA [63], CBST [64], IAST [65], and DCA [66].
In this task, Rural and Urban serve as the source and target domains, respectively. The Rural domain contains 1,296 labeled training images, while the Urban domain includes 1,366 unlabeled training images. Additionally, 677 labeled testing images from the Urban domain were used to evaluate the performance of each model. The experimental results are presented in Table V.
From the results, AdaptSegNet shows good performance in the ”building” category, but overall, it performs weakly with an mIoU of only 32.68. Its performance is particularly poor in complex scenes, such as the ”barren” class, with an IoU of only 13.62, indicating limited generalization capability. FADA achieves an overall mIoU of 31.41, highlighting its limitations in processing unlabeled target domain data, particularly in segmentation tasks involving complex terrains and vegetation. The CLAN method exhibits relatively stable performance across multiple categories with an mIoU of 33.11; however, it underperforms in the ”Forest” category, suggesting that its generalization in natural scenes requires improvement.
PyCDA demonstrates strong cross-domain segmentation capabilities with an mIoU of 36.25, excelling particularly in the ”Forest” and ”Road” categories, showcasing adaptability in diverse scenarios. CBST stands out with an mIoU of 41.32, indicating excellent performance across multiple categories, especially in ”Road” and ”Building.” IAST achieves an IoU of 86.01 in the ”Water” category, the highest among all comparative methods, resulting in an overall mIoU of 40.48 and demonstrating strong performance across various categories. DCA achieves an mIoU of 46.36, reflecting its robust cross-domain segmentation ability.
The proposed SiamSeg method outperforms all other methods with an impressive mIoU of 51.72, particularly excelling in the ”Building,” ”Road,” and ”Forest” categories, showcasing exceptional cross-domain generalization capability. Compared to other methods, SiamSeg integrates contrastive learning into self-supervised training, offering a novel approach for unsupervised domain adaptation (UDA) in semantic segmentation, and significantly enhancing the model’s performance in the target domain, achieving highly competitive results.
TABLE V: ross-domain RS image semantic segmentation results from LoveDA rural to urban. The best and second-best results are highlighted in bold and underlined, respectively, in each column. The evaluation metrics used are IoU and F1-score, where F1-score is abbreviated as F1. All values are presented as percentages (%), with larger values indicating better performance. The last column provides the average scores across all categories. Note that Segformer was trained only on the source domain and then tested directly on the target domain. ”C.L.” in the table refers to contrastive learning as proposed in this study.
Agricultural Forest Barren Water Road Building Background Overall
Method IoU IoU IoU IoU IoU IoU IoU mIoU
AdaptSegNet [29] 22.05 28.7 13.62 81.95 15.61 23.73 42.35 32.68
FADA [61] 24.79 32.76 12.7 80.37 12.76 12.62 43.89 31.41
CLAN [62] 25.8 30.44 13.71 79.25 13.75 25.42 43.41 33.11
PyCDA [63] 11.39 40.39 7.11 74.87 45.51 35.86 38.04 36.25
CBST [64] 30.05 29.69 19.18 80.05 35.79 46.1 48.37 41.32
IAST [65] 36.5 31.77 20.29 86.01 28.73 31.51 48.57 40.48
DCA [66] 36.92 42.93 16.7 80.88 51.65 49.6 45.82 46.36
SiamSeg w./o. C.L. 49.23 missing42.44 41.73 66.33 50.65 51.2 36.55 48.3
SiamSeg 49.1 47.3 47.44 71.26 56.67 52.58 37.65 51.72
IV-D2 Visualization Results
From Tables I, II, III, IV and V , we present the experimental results of SiamSeg across five cross-domain remote sensing image semantic segmentation tasks. In this section, we further visualize the model’s predictions to validate the outstanding performance of SiamSeg. Additionally, to explore the effectiveness of Contrastive Learning (C.L.), we visualize and compare the predictions of SiamSeg without C.L. (SiamSeg w.o. C.L.).
As shown in Fig. 4, we display the visualization results for the Potsdam and Vaihingen datasets, while Fig. 5 illustrates the results for the LoveDA Rural to Urban task. A comparison of these visualizations reveals that, overall, both models perform well. However, in certain local details, such as boundary recognition of buildings and roads and the smoothness of class differentiation, SiamSeg demonstrates a clear advantage over the model that does not incorporate contrastive learning. This indicates that the introduction of C.L. enhances the model’s understanding of image semantics and improves the extraction of network features.
In Fig. 4, we observe that in the task from Potsdam IRRG to Vaihingen IRRG, SiamSeg excels in recognizing the ”Building” and ”Tree” categories, while SiamSeg w.o. C.L. performs relatively poorly. This finding aligns with the experimental results presented in Table I. In the Potsdam RGB to Vaihingen IRRG task, the predictions on the left indicate that SiamSeg performs exceptionally well in recognizing ”Impervious Surface” and ”Low Vegetation,” particularly demonstrating high accuracy in predicting ”Impervious Surface” and its boundaries, consistent with the results in Table II.
Furthermore, in the task from Vaihingen IRRG to Potsdam IRRG, SiamSeg achieves outstanding performance in the ”Tree” and ”Low Vegetation” categories, accurately identifying small road structures, especially the roads within buildings in the left image. This observation corresponds with the experimental results in Table III. Although SiamSeg’s performance in the ”Clutter” category still shows some limitations, it exhibits significant improvement over the model without C.L. in this challenging category, further validating the effectiveness of the proposed contrastive learning approach. In the Vaihingen IRRG to Potsdam RGB task, SiamSeg excels in recognizing the ”Impervious Surface” category and demonstrates accurate identification in the ”Clutter” category, a finding that closely matches the experimental results in Table IV.
Observing Fig. 5, it is evident that SiamSeg generates precise predictions in cross-domain remote sensing image semantic segmentation tasks, particularly in the ”Building” and ”Barren” categories, where it delineates clear semantic boundaries. Notably, some regions classified as ”Road” in the predictions are labeled as ”Background” in the ground truth (GT). However, we do not consider this a misclassification by the model. Upon visual inspection, we find that these predicted ”Road” areas, although marked as ”Background” in the ground truth labels, exhibit characteristics typical of ”Road” from a semantic understanding perspective. This suggests that SiamSeg leverages more than just a data-driven learning mechanism; it demonstrates a stronger categorical understanding through deep feature induction and extraction, resulting in semantically plausible predictions in certain scenarios.
This phenomenon further corroborates the exceptional generalization capability of SiamSeg in handling cross-domain remote sensing image tasks, as well as the effectiveness of the contrastive learning method in complex scenarios.
Figure 4: Visualization of results on Potsdam and Vaihingen datasets. For each task, we provide 3 cases containing 4 images. From left to right, the 4 images are respectively target images, prediction of SiamSeg without contrastive learning, prediction of SiamSeg and ground truth. The cross-domain tasks from top to bottom are Potsdam IRRG to Vaihingen IRRG, Potsdam RGB to Vaihingen IRRG, Vaihingen IRRG to Potsdam IRRG and Vaihingen IRRG to Potsdam RGB. By looking at the visualization results, we can find that SiamSeg with Contrastive learning has better segmentation performance. The categories represented by the different colors are listed at the bottom of the picture with their names and colors.
Figure 5: Visualization of results on LoveDA datasets. We conduct one task which is Rural to Urban. We provide the visualization results on LoveDA dataset. Since images in the testing dataset do not have annotations, we display the results of images in the validation dataset. For this task we provide 1 case containing 4 images. From left to right, the 4 images are respectively target images, prediction of SiamSeg without contrastive learning, prediction of SiamSeg and ground truth. The categories represented by the different colors are listed at the bottom of the picture with their names and colors.
IV-D3 Ablation Studies
Effectiveness of Contrastive Learning
Using the performance of SiamSeg and SiamSeg without Contrastive Learning (C.L.) in the task from Potsdam IRRG to Vaihingen IRRG, as shown in Table I, we can clearly observe that SiamSeg outperforms SiamSeg w.o. C.L. across all categories after the introduction of C.L. This indicates that C.L. effectively enhances the model’s ability to perceive features across different categories. The additional supervisory signals provided by C.L. enable the model to learn image features more profoundly, significantly improving performance in cross-domain semantic segmentation tasks. The introduction of C.L. enriches the feature representation of the model, thereby enhancing its generalization capabilities and category differentiation.
Choice of Sim. Aug. Methods in Contrastive Learning
Within the Contrastive Learning framework, we further investigated the impact of various data augmentation methods on model performance, as shown in Table VI, including Color Jitter, Resize, and Flip. The experimental results indicate that Resize and Flip have a minimal effect on model performance, showing almost no significant differences. In contrast, the inclusion of Color Jitter resulted in a noticeable improvement in model performance. This may be attributed to the fact that Resize and Flip do not substantially alter the overall distribution of the images, preventing Contrastive Learning from extracting valuable representation information. In contrast, Color Jitter modifies the color features of the images, disrupting their original distribution, which allows Contrastive Learning to better learn useful representation information from similar images. Even without employing further data augmentation techniques such as Resize and Flip, Color Jitter alone significantly enhances model performance.
TABLE VI: The effect of different augmentation methods, such as Collor Jitter, Gray Scale, Gaussian Blur, Resize and Flip on the performance of the model. We use the average IoU and average F1-score for evaluation, where all values are in % and higher values indicate better performance. SiamSeg w. Resize denotes the version of Siamseg with the Resize augment applied in contrastive learning. C.J. denotes Collor Jitter, G.S denotes Gray Scale, G.B. denotes Gaussian Blur. The task chosen in the table is POT IRRG → VAI IRRG (Potsdam IRRG to Vaihingen IRRG).
Method mIoU mF1 Performance
SiamSeg w. Resize 63.28 74.65 ×
SiamSeg w. Flip 63.01 74.33 ×
SiamSeg w. C.J. 65.21 76.16
SiamSeg w. All 65.55 77.07
IV-D4 Computational Complexity Analysis
In this study, we employed the MixTransformers architecture from Segformer [18] as our backbone network, which has a parameter count of 85.15M. This represents an approximately 30% increase compared to DeepLabV3+ [57], which has 62.7M parameters; however, this increase is accompanied by a significant improvement in performance. Although transformer-based backbone networks typically have a higher parameter count than traditional CNN-based networks, they offer the advantage of not requiring additional auxiliary networks, such as the complex architectures needed in adversarial learning methods.
The proposed method integrates Contrastive Learning with a shared-weight backbone, adding only two MLP heads (specifically, a projection head and a prediction head). Consequently, the increase in parameter count from this addition is minimal and does not impose a significant burden on overall computational complexity. In contrast, DualGAN [58] utilizes two generators and two discriminators, with each generator having 41.28M parameters and each discriminator 6.98M, resulting in a total parameter count of 97.56M—significantly higher than the proposed SiamSeg method.
Therefore, despite the slight increase in parameter count of the transformer-based SiamSeg, its efficient weight-sharing mechanism and lightweight MLP head design ensure that its overall computational complexity remains acceptable. Moreover, compared to complex architectures like DualGAN, it achieves a significant reduction in parameter count, demonstrating better computational efficiency.
V Discussion
V-A Limitations
While the method proposed in this paper effectively addresses the performance degradation caused by unlabeled target domain data, its primary limitation lies in its reliance on joint training with labeled source domain data and unlabeled target domain data. This necessitates the separate training of a model for each cross-domain task, increasing training costs. Furthermore, due to the dependency on source domain data, training the domain adaptation model can be time-consuming, especially when computational resources are limited. When the volume of source domain data is substantial, retraining a model for each new task becomes impractical. Nevertheless, it is important to emphasize that the proposed method has a minimal dependence on target domain labels, providing a significant advantage in unsupervised scenarios. Compared to traditional methods that rely on target domain labels, our approach demonstrates greater adaptability in unlabeled environments.
V-B Future Works
Future research will further explore Source-Free Domain Adaptation (SFDA) methods, particularly in scenarios where source domain data is unavailable, to achieve more effective migration without source domain data. SFDA methods eliminate the need to reuse source domain data for training in each task, aligning more closely with practical application requirements, especially in cases where source domain data is difficult to obtain or usage is restricted. Additionally, we will focus on enhancing the model’s generalization ability, enabling it to adapt efficiently across different cross-domain tasks, thereby further reducing reliance on source domain data, lowering training costs, and increasing application flexibility.
VI Conclusion
This paper presents the SiamSeg method, which effectively addresses domain migration issues in remote sensing image cross-domain semantic segmentation tasks by integrating contrastive learning. This approach not only enhances the model’s perceptual capability for target domain features through unsupervised learning but also significantly improves the model’s cross-domain generalization ability, particularly excelling in recognizing complex categories such as buildings and roads. Experimental results demonstrate that SiamSeg achieves higher mean Intersection over Union (mIoU) and accuracy compared to existing methods, while maintaining a low computational complexity. Future work will continue to focus on reducing dependence on source domain data and exploring more efficient domain adaptation techniques to further enhance the practical application value of the model.